OpenStudy (anonymous):

Solve using transformations. 13 + m = -96 The answer is -109, is that correct?

OpenStudy (anonymous):

any body able to help?

OpenStudy (anonymous):

Transformations? Transpositions, I believe, is the correct term here. You transpose the 13 to the other side of the equation by subtraction. "Change sides, change signs."

OpenStudy (anonymous):

actually this is just a problem i found on this site that was similar to mine. Mine is: Solve using transformation: n - (-16)= 75 the term used is definitely transformation

OpenStudy (anonymous):

You are right. 13+m=-96 13+m-13=-96-13 m=-109

OpenStudy (anonymous):

that seems really simple to me. How is transformation being used here?

OpenStudy (anonymous):

If that's the answer then it's a basic algebraic equation. I don't understand how transformation has anything to do with it.

OpenStudy (anonymous):

Same here. I think someone mixed up 'transformation' with 'transposition.' I suppose it's a kind of transformation, but 'transformation' in mathematics has other more specific meanings.

OpenStudy (anonymous):

I don't think my book used the wrong word. Maybe we just dont understand the they'r asking for? In all the reading ive been doing, transformation usually involves graphing of some kind. But my book is not asking for a graph.

OpenStudy (anonymous):

In the example "Solve using transformation: n - (-16)= 75" The transformation it might be talking about could be transforming n-(-16) to n+16, but that's just multiplication. It's similar to a reflection transformation by multiplying by a negative to reflect a point over a particular axis or other line, but the term 'transformation' is really not needed here.

OpenStudy (anonymous):

ok thanks